- 2e20 out of 6.6e20 delivered.
- DI filter, replace again. Change Tuesday. Also check other filters
- replace velocity meter near target.
- examine why fan in the decay pipe is not varying smoothly. solid state controller? if so, then order a mechanical one.
- replaced humidifier last week.
- running low before outage
- 2+ hours collimator study
- Cavity down from water leak (2 hours of work, try to do with BNB work)
- Down on wednesday for some cryo pump ODH test.
- need to organize this to do every summer shutdown, not during beam time.
- Daylight savings coming.